OpenStudy (anonymous):

what multiplies to -18 and adds to -1

OpenStudy (zpupster):

xy=-18 x+y=-1 so subtract y both sides 2nd eq x+y-y=-1-y x=-1-y we now know what x is plug it in first eq (-1-y)y = -18 -y +y^2= -18 now add 18 to both sides y^2 - y +18 = 0 this can not be factored use quadratic equation the calculator gives imaginary numbers

OpenStudy (unklerhaukus):

Careful, (-1-y)y = -18 -y -y^2= -18 now add 18 to both sides -y^2 - y +18 = 0 y^2 + y - 18 = 0 with this quadratic equation there are two (real, but irrational) soultions

OpenStudy (zpupster):

thanks UnkleRhaukus